# Large Language Models Might Not Care What You Are Saying: Prompt Format Beats Descriptions

### v6

@@ Abstract
With the help of in-context learning (ICL), large language models (LLMs) have achieved impressive performance across various tasks. However, the function of descriptive instructions during ICL remains under-explored. In this work, we propose an ensemble prompt framework to describe the selection criteria of multiple in-context examples, and preliminary experiments on machine translation (MT) across six translation directions confirm that this framework boosts ICL performance. But to our surprise, LLMs might not care what the descriptions actually say, and the performance gain is primarily caused by the ensemble format, since it could lead to improvement even with random descriptive nouns. We further apply this new ensemble framework on a range of commonsense, math, logical reasoning and hallucination tasks with three LLMs and achieve promising results, suggesting again that designing a proper prompt format would be much more effective and efficient than paying effort into specific descriptions. Our code is available at https://github.com/JamyDon/Format-Beats-Descriptions.
@@
[…]
@@ 1 Introduction
The selection of in-context examples has been proven significant to the performance of ICL Rubin et al. (2022) and there have been various works on in-context example selection Agrawal et al. (2023); Li et al. (2023); Ye et al. (2023). Besides diverse approaches of selecting examples, no existing work has tried to explicitly tell LLMs in what way those specific examples are selected. We hypothesize that if LLMs are prompted with instructions describing the properties of selected in-context examples, they might learn better from these examples, since instruction following is one of LLMs’ most important qualities nowadays Ouyang et al. (2022); Peng et al. (2023); Zhang et al. (2024).
Recently, Tang et al. (2024) prompt LLMs with examples selected based on both word-level and syntax-level criteria for machine translation (MT) for better ICL performance. This inspires us to tell LLMs where different in-context examples come from when they are selected by multiple methods.
@@
There are a few studies very related to our work. Min et al. (2022) find that the labels of in-context examples do not need to be correct for classification tasks. Wei et al. (2023) find that larger language models successfully learn from in-context examples even when the labels are flipped or semantically-unrelated. Our work is different from the above in that we focus on the meaning of descriptions rather than labels in ICL and our finding is that the format of prompts is more important than carefully designed descriptions.
@@ 2.2 A New Ensemble Prompt Framework
Then, we add example-level descriptions for examples from different selection approaches and explicitly instruct the LLM to translate the test input. This prompt framework is referred to as Ensemble and is shown in Figure 1 as presented in Section 1. "Noun $A$" and "Noun $B$" describe the examples from selection $A$ and $B$ respectively. For example, the two nouns can be "words" and "syntax" to properly describe examples selected by BM25 and Polynomial respectively. In this way, we can conveniently control the example-level descriptions to tell the LLM why those examples are used.
Figure 3: Main results on XGLM and Alpaca, showing the performance gain of different prompts over the vanilla prompt, averaged over all six translation directions. Each cluster presents the results of a selection of in-context examples and each bar in it presents the result of a prompt. "Ens.", "W.", "Syn.", "Sem.", "Diff.", "Rand.", "Poly." refer to "Ensemble", "Word", "Syntax", "Semantics", "Different", "Random", "Polynomial" respectively.
@@ 2.3.4 Example Selection
To maintain consistency, all our MT experiments use 4 in-context examples. We evaluate different ways of selecting examples for comparison. Note that if all 4 examples are selected by the same method, the first two are considered examples from $A$ and the last two are considered from $B$ in the Ensemble template in Figure 1.
Random: The 4 examples are randomly sampled from the example database. We report the average result of three different random seeds.
@@ 2.3.5 Prompts
Vanilla: The normal prompt without any example-level descriptions shown in Figure 2.
Ensemble (Word + Syntax): The Ensemble prompt shown in Figure 1, with Noun $A$ and Noun $B$ being "words" and "syntax" respectively, which semantically corresponds to BM25 + Polynomial examples but is converse to Polynomial + BM25.
Ensemble (Syntax + Word): Noun $A$ and Noun $B$ are "syntax" and "words" respectively, which semantically matches Polynomial + BM25 examples but mismatches BM25 + Polynomial.
Different Ensemble (Word + Syntax): Noun $A$ and Noun $B$ are still "words" and "syntax" respectively but the qualifier "similar" is replaced with "different". This prompt aims to find out whether LLMs know the meaning of "different/similar" and care the semantics of descriptions.
Ensemble (Word + Semantics): Noun $A$ and Noun $B$ are "words" and "semantics" respectively.
Ensemble (Random + Random): For each input, Noun $A$ and Noun $B$ are different random English nouns sampled using Wonderwords [^2] https://github.com/mrmaxguns/wonderwordsmodule, aiming to explore LLMs’ understanding of descriptions.
Detailed templates and examples of the above prompts are shown in Appendix A.1.
@@ 2.4 Main Results
Selection XGLM Alpaca
@@
Table 2: Results of Vanilla baselines.
Main results are shown in Figure 3. For convenient comparison, we present the performance gain of different Ensemble prompts over Vanilla with different selections of in-context examples and the results are averaged over six translation directions.
As can be seen from the results, those "correct" prompts, exactly corresponding to the selection of in-context examples (e.g., Ensemble (Word + Syntax) with BM25 + Polynomial examples and Ensemble (Syntax + Word) with Polynomial + BM25 examples), do bring some help as expected. However, when the prompt does not correspond to the selection of examples (is "incorrect"), the performance improves as well and sometimes even more than those "correct" cases. For example, on XGLM with BM25 + Polynomial examples, Ensemble (Syntax + Word) improves more than Ensemble (Word + Syntax), even though the former is completely reversed. On Alpaca with BM25 + Polynomial examples, Ensemble (Word + Semantics) improves more than Ensemble (Word + Syntax), albeit the examples with similar syntax do not necessarily bear similar semantics. More interestingly, Different Ensemble (Word + Syntax), telling the LLM that the in-context examples are with different properties, is able to beat "correct" prompts sometimes (e.g., on XGLM with BM25 + Polynomial examples and Alpaca with Polynomial + BM25 examples).
@@ 2.5 Ablation Study
Detailed templates and examples of the above prompts are shown in Appendix A.2.
Figure 4: Ablation studies over the organization of the prompt, showing the performance gain of different prompts over Vanilla, averaged over all six translation directions. "Rand.", "Poly.", "Ens.", "Sgl.", "V.", "Trans." refer to "Random", "Polynomial", "Ensemble", "Single", "Vanilla", "Translate" respectively.
@@
Results are presented in Figure 4, showing that removing one or two example-level descriptions or removing the random noun describing the property of in-context examples hurt the performance gain in most cases. On XGLM, only the original Ensemble format performs better than Vanilla. On Alpaca, Single (Random), Single (Example) and Vanilla (Translate) bring less improvement than Ensemble (Random + Random) in more than half of the cases.
@@ 2.6 Analysis via Attention Weights
To have a better idea of the internal mechanism of LLMs when prompted with different prompts, we calculate the attention weights between different prompt components. We focus on three components: in-context examples (from $A$ or $B$, denoted by "Example-A" and "Example-B"), the target position (denoted by "Target") where the model starts to generate predictions (following Wang et al. Wang et al. (2023), we use the final token in the input) and the two descriptive nouns ("Noun-A" and "Noun-B"). We obtain the attention weights averaged over all attention heads from the attention matrix across all the layers. All the results are averaged over all six language directions.
Results comparing Ensemble (Word + Syntax) (EWS) and Ensemble (Random + Random) (ERR) on XGLM with BM25 + Polynomial examples are presented in Figure 5 (for results on Alpaca, refer to Appendix B). If the model really cares what the descriptions say, its attention to meaningful descriptive nouns (in EWS) should be much greater than those meaningless (in ERR). However, in most cases, EWS performs no higher than ERR, indicating that the model does not really care what the descriptive nouns actually are. "Target to Noun-A" is a special case, where EWS is high in shallow layers. But in deeper layers, EWS falls behind and ERR takes the lead. This shows that the model might pay more attention to the meaningful noun when understanding the context in shallow layers but gradually forgets it when it comes to generation in deeper layers. In a word, the attention weights further confirm our claim that LLMs do not really care what the descriptive nouns are in most cases.
@@ 3.1.1 Datasets
The number of test inputs for each dataset is listed in Table 3. Details of splitting training set (example database) and test set are in Appendix C.
@@ 3.1.3 Language Models
We evaluate three open source LLMs with around 7B parameters, including Alpaca Taori et al. (2023), Llama3 Grattafiori et al. (2024), and Mistral Jiang et al. (2023). To assess the effect of ERR on more powerful models, we also evaluate GPT-3.5 Ouyang et al. (2022). We use Llama-3.1-8B, Mistral-7B-Instruct-v0.2 and gpt-3.5-turbo-0125333https://openai.com/api/ for Llama3, Mistral and GPT-3.5 respectively.
@@ 3.1.5 Prompts
We compare ERR with Vanilla across different datasets and LLMs. Given that these tasks usually involve reasoning, on which chain-of-thought (CoT) is commonly utilized Wei et al. (2022), we experiment both without CoT ("w/o CoT", which are identical to the original templates) and with CoT ("w/ CoT"). This allows us to examine both the orthogonality and compatibility with CoT of ERR , as well as assess its performance across various models and tasks. Specifically, we evaluate Vanilla (w/o CoT), Vanilla (w/ CoT), ERR (w/o CoT), and ERR (w/ CoT). Due to space constraints, examples of prompt templates discussed in this section are provided in Appendix D.
@@ 3.2 Results of Small-scale Models
Results across all nine datasets and three small-scale models (Alpaca, Llama3 and Mistral) are illustrated in Figure 6.
The results demonstrate that ERR (w/ CoT), achieved by integrating CoT with our proposed prompt framework, either significantly outperforms or matches Vanilla (w/ CoT) in 25 out of 27 experiments (covering 9 datasets and 3 models). The exceptions are Alpaca on the Sports dataset and Mistral on the AQuA dataset, where ERR (w/ CoT) shows somewhat lower performance compared to Vanilla (w/ CoT). When CoT is not incorporated, ERR generally performs much better than or on par with Vanilla, except for the Sports dataset with Llama3, where ERR performs a little poorer.
@@
In summary, without using any carefully designed selection methods, directly filling the randomly selected examples into the ERR framework brings significant improvement to various reasoning tasks and even alleviates the hallucination of models in most cases, no matter how meaningless and incorrect the example-level descriptions are. Moreover, ERR can work perfectly with CoT. Therefore, at least for relatively small models, this simple but effective trick is worth introducing into prompt engineering for various tasks.
We also experiment with Llama2 Touvron et al. (2023b) and the results are in Appendix E. The overall trend is consistent with Llama3.
@@ 3.4 Discussion
Based on the experiments conducted on both small-scale and large-scale models, we can conclude that ERR is a simple yet practical and universal prompt framework. It can enhance problem-solving capabilities in small models and be applied to large models without the risk of performance degradation due to the meaningless noise within it. In other words, there might be less need to meticulously select examples or design detailed descriptions. Instead, you can uniformly and efficiently apply ERR to various tasks with different models.
@@ 4 Related Work
Rubin et al. Rubin et al. (2022) suggest that LLMs’ ICL performance strongly depends on the selection of in-context examples. In consequence, many works have been trying to explore ways of selecting better in-context examples in recent years. Li et al. Li et al. (2023) train a unified in-context example retriever across a wide range of tasks. Ye et al. Ye et al. (2023) select examples based on both relevance and diversity, with the help of determinantal point processes. Agrawal et al. Agrawal et al. (2023) ensure n-gram coverage to select better examples for MT. Kumar et al. Kumar et al. (2023) train an in-context example scorer for MT based on several features. Tang et al. Tang et al. (2024) propose to combine both word-level and syntax-level coverage when selecting examples for MT.
@@ 5 Conclusion
Appendix A Prompt Templates of MT Experiments
Figure 8: Template and example of Vanilla.
Figure 9: Template and example of Ensemble (Word + Syntax).
Figure 10: Template and example of Ensemble (Syntax + Word).
Figure 11: Template and example of Different Ensemble (Word + Syntax).
Figure 12: Template and example of Ensemble (Word + Semantics).
@@ Appendix A Prompt Templates of MT Experiments
A.1 Main Experiments
Templates and prompt examples of Vanilla, Ensemble (Word + Syntax), Ensemble (Syntax + Word), Different Ensemble (Word + Syntax), Ensemble (Word + Semantics) and Ensemble (Random + Random) are shown in Figure 8-13.
A.2 Ablations
@@
Templates and prompt examples of Ensemble (Random + Random), Single (Random), Single (Example), Vanilla (Translate) are shown in Figure 13-16.
Appendix B Attention Weights on Alpaca

### v7

@@ Abstract
With the help of in-context learning (ICL), large language models (LLMs) have achieved impressive performance across various tasks. However, the function of descriptive instructions during ICL remains under-explored. In this work, we propose an ensemble prompt framework to describe the selection criteria of multiple in-context examples, and preliminary experiments on machine translation (MT) across six translation directions confirm that this framework boosts ICL performance. But to our surprise, LLMs might not care what the descriptions actually say, and the performance gain is primarily caused by the ensemble format, since it could lead to improvement even with random descriptive nouns. We further apply this new ensemble framework on a range of commonsense, math, logical reasoning and hallucination tasks with three LLMs and achieve promising results, suggesting again that designing a proper prompt format would be much more effective and efficient than paying effort into specific descriptions.
@@
$\dagger$$\dagger$}\$footnotetext:  These authors contributed equally.
@@ 1 Introduction
The selection of in-context examples has been proven significant to the performance of ICL Rubin et al. (2022) and there have been various works on in-context example selection Agrawal et al. (2023); Li et al. (2023); Ye et al. (2023). Besides diverse approaches of selecting examples, no existing work has tried to explicitly tell LLMs in what way those specific examples are selected. We hypothesize that if LLMs are prompted with instructions describing the properties of selected in-context examples, they might learn better from these examples, since instruction following is one of LLMs’ most important qualities nowadays Ouyang et al. (2022); Peng et al. (2023); Zhang et al. (2024). Tang et al. (2024) prompt LLMs with examples selected based on both word-level and syntax-level criteria for machine translation (MT) for better ICL performance. This inspires us to tell LLMs where different in-context examples come from when they are selected by multiple methods.
@@
There are a few studies very related to our work. Min et al. (2022) find that the labels of in-context examples do not need to be correct for classification tasks. Wei et al. (2023) find that larger language models learn from in-context examples even when the labels are flipped or unrelated. Srivastava et al. (2024) demonstrate that optimizing examples is less effective in some tasks given a high-quality task instruction. Our work is different from the above in that we focus on the meaning of descriptions rather than labels or examples in ICL and our finding is that the format of prompts is more important than carefully designed descriptions.
@@ 2.2 A New Ensemble Prompt Framework
Then, we add example-level descriptions for examples from different selection approaches and explicitly instruct the LLM to translate the test input. This prompt framework is referred to as Ensemble and is shown in Figure 1 as presented in Section 1. "Noun AA" and "Noun BB" describe the examples from selection $A$ and $B$ respectively. For example, the two nouns can be "words" and "syntax" to properly describe examples selected by BM25 and Polynomial respectively. In this way, we can conveniently control the example-level descriptions to tell the LLM why those examples are used.
@@ 2.3.4 Example Selection
To maintain consistency, all our MT experiments use four in-context examples (4-shot). We evaluate different ways of selecting examples for comparison. Note that if all four examples are selected by the same method, the first two are considered examples from $A$ and the last two are considered from $B$ in the Ensemble template in Figure 1.
Random: The four examples are randomly sampled from the example database. We report the average result of three different random seeds.
@@ 2.3.5 Prompts
Vanilla: The normal prompt without any example-level descriptions as shown in Figure 2.
Ensemble (Word + Syntax): Shown in Figure 3(a), Noun $A$ and Noun $B$ are "words" and "syntax" respectively, which semantically corresponds to BM25 + Polynomial examples but is converse to Polynomial + BM25.
Ensemble (Syntax + Word): Shown in Figure 3(b), Noun $A$ and Noun $B$ are "syntax" and "words" respectively, which semantically matches Polynomial + BM25 examples but mismatches BM25 + Polynomial.
Different Ensemble (Word + Syntax): Shown in Figure 3(c), Noun $A$ and Noun $B$ are still "words" and "syntax" respectively but the qualifier "similar" is replaced with "different". This aims to find out whether LLMs pay attention to the meaning of "different/similar" and care the semantics of descriptions.
Ensemble (Word + Semantics): Shown in Figure 3(d), Noun $A$ and Noun $B$ are "words" and "semantics" respectively, which does not semantically match any of our example selection methods.
Ensemble (Random + Random): Shown in Figure 3(e), for each input, Noun $A$ and Noun $B$ are different random English nouns sampled using Wonderwords  [^2] https://github.com/mrmaxguns/wonderwordsmodule, aiming to explore LLMs’ understanding of descriptions.
(a) Ensemble (Word + Syntax).
(b) Ensemble (Syntax + Word).
(c) Different Ensemble (Word + Syntax).
(d) Ensemble (Word + Semantics).
(e) Ensemble (Random + Random).
Figure 3: Templates and Alpaca’s examples of Ensemble prompts.
Figure 4: Main results on XGLM and Alpaca, showing the performance gain of different prompts over the Vanilla prompt, averaged over all six translation directions. Each cluster presents the results of a selection of in-context examples and each bar in it presents the result of a prompt. "Ens.", "W.", "Syn.", "Sem.", "Diff.", "Rand.", "Poly." refer to "Ensemble", "Word", "Syntax", "Semantics", "Different", "Random", "Polynomial", respectively.
@@ 2.4 Main Results
Example Selection XGLM Alpaca
@@
Table 2: Results of Vanilla baselines of XGLM and Alpaca with different example selection methods, averaged over six translation directions.
Main results are shown in Figure 4. For convenient comparison, we present the performance gain of different Ensemble prompts over Vanilla with different selections of in-context examples and the results are averaged over six translation directions. For detailed results of different translation directions, please refer to Appendix B.
As can be seen from the results, those "correct" prompts, exactly corresponding to the selection of in-context examples (e.g., Ensemble (Word + Syntax) with BM25 + Polynomial examples and Ensemble (Syntax + Word) with Polynomial + BM25 examples), do bring some help as expected. However, when the prompt does not correspond to the selection of examples (i.e., is "incorrect"), the performance improves as well and sometimes even more than those "correct" cases. For example, on XGLM with BM25 + Polynomial examples, Ensemble (Syntax + Word) improves more than Ensemble (Word + Syntax), even though the former is completely reversed. On Alpaca with BM25 + Polynomial examples, Ensemble (Word + Semantics) improves more than Ensemble (Word + Syntax), albeit the examples with similar syntax do not necessarily bear similar semantics. More interestingly, Different Ensemble (Word + Syntax), telling the LLM that the in-context examples are with different properties, is able to beat "correct" prompts sometimes (e.g., on XGLM with BM25 + Polynomial examples and Alpaca with Polynomial + BM25 examples).
@@ 2.5 Ablation Study
Detailed templates and examples of the above prompts are presented in Appendix A.
Figure 5: Ablation studies over the organization of the prompt, showing the performance gain of different prompts over Vanilla, averaged over all six translation directions. "Rand.", "Poly.", "Ens.", "Sgl.", "V.", "Trans." refer to "Random", "Polynomial", "Ensemble", "Single", "Vanilla", "Translate", respectively.
@@
Results are presented in Figure 5, showing that removing one or two example-level descriptions or removing the random noun describing the property of in-context examples hurt the performance gain in most cases. On XGLM, only the original Ensemble format performs better than Vanilla. Alpaca exhibits an abnormal trend when prompted with Polynomial and BM25 + Polynomial examples, where Ensemble (Random + Random) cannot outperform other prompts. This may be due to that Alpaca is instruction-tuned and the Single or Vanilla (Translate) prompts are also friendly to it in some cases because of the post-training stage. But overall, Single (Random), Single (Example) and Vanilla (Translate) still bring less improvement than Ensemble (Random + Random) in more than half of the cases.
@@ 2.6 Analysis via Attention Weights
To have a better idea of the internal mechanism of LLMs when prompted with different prompts, we compute the attention weights between different prompt components. We focus on three components: in-context examples (from $A$ or $B$, denoted by "Example-A" and "Example-B"), the target position (denoted by "Target") where the model starts to generate predictions (following Wang et al. Wang et al. (2023), we use the final token in the input) and the two descriptive nouns ("Noun-A" and "Noun-B"). We obtain the attention weights averaged over all attention heads from the attention matrix across all the layers. All the results are averaged over all six language directions.
Results comparing Ensemble (Word + Syntax) (EWS) and Ensemble (Random + Random) (ERR) on XGLM with BM25 + Polynomial examples are presented in Figure 6 (for results on Alpaca, refer to Appendix C). If the model really cares what the descriptions say, its attention to meaningful descriptive nouns (in EWS) should be much greater than those meaningless (in ERR). However, in most cases, EWS performs no higher than ERR, indicating that the model does not really care what the descriptive nouns actually are. "Target to Noun-A" is a special case, where EWS is high in shallow layers. But in deeper layers, EWS falls behind and ERR takes the lead. This shows that the model might pay more attention to the meaningful noun when understanding the context in shallow layers but gradually forgets it when it comes to generation in deeper layers. In a word, the attention weights further confirm our claim that LLMs do not really care what the descriptive nouns are in most cases.
@@ 3.1.1 Datasets
The number of test inputs for each dataset is listed in Table 3. Details of splitting training set (example database) and test set are in Appendix D.
@@ 3.1.3 Language Models
We experiment with both instruction-tuned and non-instruction-tuned models to see whether our findings could extend to different kinds of models. We evaluate three frequently used open source LLMs with around 7B parameters, including Alpaca Taori et al. (2023), Llama3 Llama Team (2024), and Mistral Jiang et al. (2023), among which Llama3 is a base model before instruction tuning. To assess the effect of ERR on more powerful models, we also evaluate GPT-3.5 Ouyang et al. (2022)  [^3] We choose this model because it is a commonly used cost-effective API-based LLM and a de facto black box baseline.. We use Llama-3.1-8B, Mistral-7B-Instruct-v0.2 and gpt-3.5-turbo-0125  [^4] https://openai.com/api/ for Llama3, Mistral and GPT-3.5 respectively.
@@ 3.1.5 Prompts
We compare ERR with Vanilla across different datasets and LLMs. Given that these tasks usually involve reasoning, on which chain-of-thought (CoT) is commonly utilized Wei et al. (2022), we experiment both without CoT ("w/o CoT", which are identical to the original templates) and with CoT ("w/ CoT"). This allows us to examine both the orthogonality and compatibility with CoT of ERR , as well as assess its performance across various models and tasks. Specifically, we evaluate Vanilla (w/o CoT), Vanilla (w/ CoT), ERR (w/o CoT), and ERR (w/ CoT). Due to space constraints, examples of prompt templates discussed in this section are provided in Appendix E.
@@ 3.2 Results of Small-scale Models
Results across all nine datasets and three small-scale models (Alpaca, Llama3 and Mistral) are illustrated in Figure 7. Detailed results are presented in Appendix B.
The results demonstrate that ERR (w/ CoT), achieved by integrating CoT with our proposed prompt framework, either significantly outperforms or matches Vanilla (w/ CoT) in 25 out of 27 experiments (covering nine datasets and three models). The exceptions are Alpaca on the Sports dataset and Mistral on the AQuA dataset, where ERR (w/ CoT) shows somewhat lower performance compared to Vanilla (w/ CoT). When CoT is not incorporated, ERR generally performs much better than or on par with Vanilla, except for the Sports dataset with Llama3, where ERR performs a little poorer.
@@
In summary, without using any carefully designed selection methods, directly filling the randomly selected examples into the ERR framework brings significant improvement to various reasoning tasks and even alleviates the hallucination of models in most cases, no matter how meaningless and incorrect the example-level descriptions are. Moreover, ERR can work perfectly with CoT. Therefore, at least for relatively small models, this simple yet effective trick is worth introducing into prompt engineering for various tasks.
We also experiment with Llama2 Llama Team (2023) and the results are in Appendix F. The overall trend is consistent with Llama3.
@@ 3.4 Discussion
Based on the experiments conducted on both small-scale and large-scale models, we can conclude that ERR is a simple yet practical and universal prompt framework. It can enhance problem-solving capabilities of small models and be applied to large models without the risk of performance degradation due to the meaningless noise within it. In other words, there might be less need to meticulously select examples or design detailed descriptions. Instead, you can uniformly and efficiently apply ERR to various tasks with different models.
As analyzed in Section 2.6, the ERR framework can work because LLMs pay less attention to the descriptive nouns while being more sensitive to the overall prompt format. We conjecture that the underlying reason could be that LLMs have been presented with many patterns similar to ERR during pre-training and thus perform better when presented with ERR prompts Chen et al. (2024). However, due to lack of access to the pre-training process of LLMs (either open-source or close-source), we cannot further validate our conjecture more solidly and our understanding of the deeper mechanism remains limited to superficial analysis, which is one of the limitations of this work.
@@ 4 Related Work
Rubin et al. Rubin et al. (2022) suggest that LLMs’ ICL performance strongly depends on the selection of in-context examples. In consequence, many works have been trying to explore ways of selecting better in-context examples in recent years. Li et al. Li et al. (2023) train a unified in-context example retriever across a wide range of tasks. Ye et al. Ye et al. (2023) select examples based on both relevance and diversity, with the help of determinantal point processes. Agrawal et al. Agrawal et al. (2023) ensure n-gram coverage to select better examples for MT. Kumar et al. Kumar et al. (2023) train an in-context example scorer for MT based on several features. Tang et al. Tang et al. (2024) combine both word-level and syntax-level coverage when selecting examples for MT.
@@ 5 Conclusion
Acknowledgments
This work is supported by the National Natural Science Foundation of China (62076008).
@@ Acknowledgments
Appendix A Prompts for MT Ablation Study
@@
Templates and prompt examples of Ensemble (Random + Random), Single (Random), Single (Example), Vanilla (Translate) are shown in Figure 9-12.
Appendix B Full Experimental Results
Prompt Selection-A Selection-B Into EN Out of EN Avg.
DE FR RU DE FR RU
Vanilla Random Random 63.65 71.40 52.37 40.34 54.58 42.07 54.07
BM25 BM25 64.32 71.83 51.15 42.53 56.58 43.58 55.00
Polynomial Polynomial 64.29 71.25 53.47 43.23 55.15 45.73 55.52
BM25 Polynomial 65.34 72.01 53.74 43.41 56.48 46.04 56.17
Polynomial BM25 64.66 72.04 53.04 44.37 56.56 46.43 56.18
Ensemble (Word + Syntax) Random Random 65.07 72.41 54.14 43.09 56.25 42.07 55.50
BM25 BM25 66.02 72.81 54.13 43.77 56.81 43.19 56.12
Polynomial Polynomial 65.98 72.66 54.67 44.37 58.13 45.49 56.88
BM25 Polynomial 66.08 72.58 54.29 43.51 57.07 45.26 56.47
Polynomial BM25 65.89 73.03 54.06 43.78 57.60 46.87 56.87
Ensemble (Syntax + Word) Random Random 65.04 72.37 54.15 42.95 56.27 41.85 55.40
BM25 BM25 66.20 72.41 53.73 43.39 57.24 42.40 55.90
Polynomial Polynomial 66.16 72.55 54.66 44.75 56.82 44.38 56.55
BM25 Polynomial 65.99 72.77 54.21 44.05 57.26 45.09 56.56
Polynomial BM25 65.96 72.96 54.12 43.47 57.51 46.94 56.83
Diff. Ensemble (Word + Syntax) Random Random 65.09 72.53 53.84 42.65 56.15 41.38 55.27
BM25 BM25 65.95 72.33 54.06 43.42 57.27 41.88 55.82
Polynomial Polynomial 65.98 72.58 54.76 44.60 58.36 45.95 57.04
BM25 Polynomial 66.04 72.49 54.75 44.03 57.70 45.95 56.83
Polynomial BM25 66.17 73.17 54.10 44.17 57.48 46.02 56.85
Ensemble (Word + Semantics) Random Random 65.09 72.26 54.14 42.64 56.11 42.61 55.48
BM25 BM25 66.32 72.25 53.89 43.29 57.02 42.51 55.88
Polynomial Polynomial 66.36 72.46 54.62 44.17 57.24 45.61 56.74
BM25 Polynomial 65.80 72.74 54.40 44.06 56.86 45.54 56.57
Polynomial BM25 66.06 72.99 53.47 43.24 57.55 46.93 56.71
Ensemble (Random + Random) Random Random 65.37 72.54 54.01 42.59 56.22 40.41 55.19
BM25 BM25 66.41 72.72 54.24 43.08 56.59 41.44 55.75
Polynomial Polynomial 66.28 72.56 54.70 43.81 57.42 43.22 56.33
BM25 Polynomial 65.99 72.77 54.04 44.85 57.25 45.39 56.72
Polynomial BM25 65.73 73.26 53.73 43.27 56.78 44.73 56.25
Table 4: Full MT results of XGLM.
Prompt Selection-A Selection-B Into EN Out of EN Avg.
DE FR RU DE FR RU
Vanilla Random Random 69.88 76.46 57.80 42.52 56.61 29.25 55.42
BM25 BM25 69.08 76.41 58.52 43.65 57.34 32.63 56.27
Polynomial Polynomial 69.65 75.79 58.77 43.55 56.60 32.39 56.13
BM25 Polynomial 69.34 76.02 58.38 43.31 56.79 33.21 56.18
Polynomial BM25 69.03 76.11 57.84 42.20 55.81 31.93 55.49
Ensemble (Word + Syntax) Random Random 69.86 76.64 57.57 43.51 57.25 30.79 55.94
BM25 BM25 69.44 76.21 57.60 44.41 58.16 31.26 56.18
Polynomial Polynomial 69.86 76.06 58.26 44.46 57.06 33.27 56.50
BM25 Polynomial 69.58 76.09 57.98 43.84 57.57 32.85 56.32
Polynomial BM25 69.41 76.32 57.78 42.79 58.18 31.07 55.93
Ensemble (Syntax + Word) Random Random 69.80 76.73 57.53 43.50 57.31 30.70 55.93
BM25 BM25 69.46 76.15 57.63 44.58 58.46 32.09 56.40
Polynomial Polynomial 69.76 76.11 58.14 44.11 56.64 33.24 56.33
BM25 Polynomial 69.60 76.11 57.91 43.88 57.81 32.22 56.26
Polynomial BM25 69.64 76.11 57.77 42.42 58.43 31.41 55.96
Diff. Ensemble (Word + Syntax) Random Random 69.77 76.63 57.46 43.67 57.48 30.55 55.93
BM25 BM25 69.44 76.33 57.70 44.24 58.48 31.76 56.33
Polynomial Polynomial 69.75 76.07 58.12 44.09 57.31 32.46 56.30
BM25 Polynomial 69.54 76.21 57.68 43.61 57.49 32.27 56.13
Polynomial BM25 69.49 76.23 57.65 42.79 58.20 32.09 56.08
Ensemble (Word + Semantics) Random Random 69.88 76.70 57.51 43.30 57.35 30.98 55.96
BM25 BM25 69.44 76.20 57.65 44.47 57.88 32.48 56.35
Polynomial Polynomial 69.82 76.10 58.35 44.00 57.25 33.45 56.50
BM25 Polynomial 69.57 76.23 58.12 44.07 57.73 32.79 56.42
Polynomial BM25 69.49 76.17 57.99 43.20 58.49 31.66 56.17
Ensemble (Random + Random) Random Random 69.77 76.61 57.38 43.52 57.55 30.90 55.95
BM25 BM25 69.46 76.24 57.58 44.40 58.25 33.23 56.53
Polynomial Polynomial 69.63 75.93 57.77 44.36 56.73 33.45 56.31
BM25 Polynomial 69.55 76.01 57.86 42.75 57.80 32.63 56.10
Polynomial BM25 69.74 76.04 57.73 43.33 58.30 33.27 56.40
Table 5: Full MT results of Alpaca.
Model Template Performance
Date SgyQA CSQA Sports LF TO GSM8K AQuA KU Avg.
LLaMA-2-7B Vanilla w/ CoT 3.01 50.99 21.70 50.20 48.72 30.74 22.29 21.65 45.24 32.73
ERR w/ CoT 45.21 59.78 55.36 79.42 56.03 41.55 21.15 23.62 85.71 51.98
Vanilla w/o CoT 0.27 47.33 19.25 54.42 48.91 30.07 0.91 23.23 52.38 30.75
ERR w/o CoT 27.67 59.88 23.91 16.77 1.58 36.49 5.91 22.44 69.05 29.30
LLaMA-3.1-8B Vanilla w/ CoT 9.04 51.28 24.41 55.32 52.08 35.47 49.05 24.80 69.05 41.17
ERR w/ CoT 38.63 55.53 31.04 74.60 51.58 38.18 48.52 26.77 69.05 48.21
Vanilla w/o CoT 7.95 41.90 28.42 54.22 51.98 37.50 1.14 16.93 28.57 29.84
ERR w/o CoT 13.70 59.68 36.12 50.00 50.79 35.81 4.70 22.83 57.14 36.75
Alpaca-7B Vanilla w/ CoT 26.03 60.38 50.53 80.12 52.27 35.81 6.44 18.11 71.43 44.57
ERR w/ CoT 26.85 60.47 50.04 74.40 54.74 35.14 6.75 20.08 71.43 44.43
Vanilla w/o CoT 25.21 60.77 46.27 60.24 55.63 35.47 4.70 24.80 64.29 41.93
ERR w/o CoT 24.93 61.17 50.12 61.04 55.14 34.80 4.78 24.41 64.29 42.30
Mistral-7B Vanilla w/ CoT 6.30 57.71 50.37 61.24 48.22 31.08 43.21 29.13 73.81 44.56
ERR w/ CoT 36.99 66.60 61.83 84.94 64.03 33.78 43.97 25.98 83.33 55.72
Vanilla w/o CoT 0.27 49.70 20.97 78.21 46.05 31.08 1.06 18.90 61.90 34.24
ERR w/o CoT 13.15 64.13 55.86 77.11 61.86 49.32 7.96 22.44 66.67 46.50
Table 6: Full results of Date Understanding (Date), StrategyQA (SgyQA), CSQA, Sports Understanding (Sports), Logical Fallacy (LF), Three Objects (TO), GSM8K, AQuA and Known Unknowns (KU).
Full results of MT are presented in Table 4 and 5. Full results of other tasks in Section 3.2 are presented in Table 6.
Appendix C Attention Weights on Alpaca